OpenStudy (anonymous):

13.) The measures of two supplementary angles are 3x + 28 and 4x – 2. What are the measures of the two angles? ..............MEDALS.............

OpenStudy (anonymous):

Supplementary angles are two angles whose sum is 180 degrees. Now, you write an equation.

OpenStudy (anonymous):

3*90+28 AND 4*90-2 ?

OpenStudy (anonymous):

what is the equation needed to solve this problem?

OpenStudy (anonymous):

Read my first post as to what supplementary angles/

OpenStudy (anonymous):

idk thants y i put it on hear lol i have no clue

OpenStudy (anonymous):

do you know what the word SUM means?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

what does it mean?

OpenStudy (anonymous):

like what somthing equals

OpenStudy (anonymous):

No. Sum means to add so if you have two angles, and their sum is 180 degrees, then you take the two angles and ADD them together and set them equal to 180 degrees. Let's see you do that.

OpenStudy (anonymous):

but how do i add the angles if idk them

OpenStudy (anonymous):

Look at your problem. One angle is 3x + 28 and the other angle is 4x - 2.

OpenStudy (anonymous):

So whats your equation needed to solve this problem?

OpenStudy (anonymous):

add?

OpenStudy (anonymous):

Yes. We just said, that SUM means to ADD.

OpenStudy (anonymous):

So whats your equation?

OpenStudy (anonymous):

3x+28+4x-2

OpenStudy (anonymous):

yes, but whats your equation?

OpenStudy (anonymous):

idk addition?

OpenStudy (anonymous):

Do you know what an equation is?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

You never solved an equation in your math class?

OpenStudy (anonymous):

probobly have but im just now starting to care so i probobly missed that

OpenStudy (anonymous):

Ok. x + 5 = 8 is an equation

OpenStudy (anonymous):

You have the = sign

OpenStudy (anonymous):

So what would be your equation to solve your problem?

OpenStudy (anonymous):

so i solve for x now ?

OpenStudy (anonymous):

I want to see your equation. You cant solve for x without an equation.

OpenStudy (anonymous):

what is your equation?

OpenStudy (anonymous):

by adding ?

OpenStudy (anonymous):

Yes, we said that SUM means to add.

OpenStudy (anonymous):

so the answer is -3?

OpenStudy (anonymous):

Would you please write the equation needed to solve your problem?

OpenStudy (anonymous):

x+5=8

OpenStudy (anonymous):

Now, write the equation needed to solve the original problem you posted about the supplementary angles.

OpenStudy (anonymous):

Can you write it?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

The 2 angles are called 3x+28 and 4x - 2. Their SUM is 180 degrees. So the equation you would need is: 3x + 28 + 4x - 2 = 180 Does that make sense?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Are you sure it makes sense?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

well?